OpenStudy (anonymous):

-6x-3y=12. Find the slope and the y-intercept.

OpenStudy (mrnood):

first re-arrange the equation with all the tems with y in on one side, and all the other terms on the other side

OpenStudy (mrnood):

Post your answer and I will help you further..

OpenStudy (jmark):

Equation of line is given as y = mx+b -6x-3y = 12 -3y = 6x+12 divide the equation by -3 y = -2x-4 slope = -2 y intercept = -4 ICSE board syllabus for class 8 - http://icse.edurite.com/icse-syllabus/icse-syllabus-for-class-8.html